### v1

# MagicTime: Time-lapse Video Generation Models as Metamorphic Simulators

Abstract
Recent advances in Text-to-Video generation (T2V) have achieved remarkable success in synthesizing high-quality general videos from textual descriptions. A largely overlooked problem in T2V is that existing models have not adequately encoded physical knowledge of the real world, thus generated videos tend to have limited motion and poor variations. In this paper, we propose MagicTime, a metamorphic time-lapse video generation model, which learns real-world physics knowledge from time-lapse videos and implements metamorphic generation. First, we design a MagicAdapter scheme to decouple spatial and temporal training, encode more physical knowledge from metamorphic videos, and transform pre-trained T2V models to generate metamorphic videos. Second, we introduce a Dynamic Frames Extraction strategy to adapt to metamorphic time-lapse videos, which have a wider variation range and cover dramatic object metamorphic processes, thus embodying more physical knowledge than general videos. Finally, we introduce a Magic Text-Encoder to improve the understanding of metamorphic video prompts. Furthermore, we create a time-lapse video-text dataset called ChronoMagic, specifically curated to unlock the metamorphic video generation ability. Extensive experiments demonstrate the superiority and effectiveness of MagicTime for generating high-quality and dynamic metamorphic videos, suggesting time-lapse video generation is a promising path toward building metamorphic simulators of the physical world.  [^1] Our project aims to help reproduce Sora, supporting Open-Sora-Plan or other DiT-based T2V models.
Figure 1: The illustration of the difference between (a) general videos, and (b) metamorphic videos. Compared to general videos, metamorphic videos contain physical knowledge, long persistence, and strong variation, making them difficult to generate. The general videos are generated by the Animatediff (animatediff, ). Our MagicTime greatly improves capabilities of metamorphic video generation.
1 Introduction
Recent advances in Text-to-Video (T2V) generation models have been driven by the emergence of diffusion models. These models are typically based on either Transformer architectures (Videopoet, ; Latte, ) or U-Net architectures (VideoComposer, ; Modelscope, ; Align_your_latents, ; animatediff, ). The former approach tokenizes videos and incorporates positional encoding for training, whereas the latter extends the 2D U-Net with a temporal feature extraction module for pseudo-3D training. However, the videos generated by these models primarily exhibit camera movement, lacking a continuous subject metamorphic process and limited incorporation of physical knowledge. We categorize such videos as general videos, illustrated in Fig. 1a. Since the training datasets predominantly consist of general videos, current T2V models struggle to produce videos depicting complex phenomena like seed germination, ice melting, or flower blooming, which are characterized by minimal physical content, brief duration, and limited variation.
Compared to general videos, we have observed a category of videos that typically encompasses the subject’s entire transformation process, thus addressing the inherent limitations of the former. We term this category as metamorphic videos, which encode a more comprehensive representation of world knowledge, as illustrated in Fig. 1b. Such videos have the potential to enhance the capacity of video generation models to accurately depict the real world. Consequently, the generation of metamorphic videos is of paramount importance for a wide range of applications, despite being relatively underexplored in existing research. While the study (freebloom, ) demonstrates the ability to create similar effects by repeatedly inferring and concatenating general videos, it falls short in encoding physical knowledge and lacks generalization through a prompt strategy. In contrast, we aim to develop an end-to-end time-lapse video generation method for adaptively encoding physical knowledge with good generalization.
Time-lapse videos provide detailed documentation of an object’s complete metamorphosis, possessing the essential characteristics of metamorphic videos. Consequently, we utilize time-lapse videos to validate our hypothesis. A straightforward approach is to leverage existing powerful T2V models to directly generate metamorphic videos. Since metamorphic videos and general videos have intrinsic gaps, current methods struggle to produce satisfactory metamorphic videos directly. The core challenges include: (1) Metamorphic time-lapse videos contain more physics, requiring stronger structural extraction; (2) It has longer persistence, fully covering object metamorphosis and being sensitive to training strategies; (3) Metamorphic time-lapse videos have greater variation, potentially transforming from seed to flower, demanding high-quality data.
To tackle the challenge of generating metamorphic videos, we introduce a novel framework, MagicTime, which excels in creating video content that is Magically compressed in Time. Initially, we present the MagicAdapter to encode more physical knowledge in metamorphic videos into the feature extraction, expanding the metamorphic video generation capabilities of pre-trained T2V models without altering their core general video generation abilities. Subsequently, we propose a Dynamic Frames Extraction strategy that enables the model to adapt to the characteristics of time-lapse training videos and prioritize metamorphic video features. Moreover, we introduce a Meta Text-Encoder to refine prompt understanding and distinguish between metamorphic and general prompts. As shown in Fig. 1b, our MagicTime generates text-aligned, coherent, high-quality metamorphic videos across various styles, coordinated with carefully designed techniques.
Table 1: Comparison of the characteristics of our ChronoMagic benchmark with existing time-lapse datasets. We compare time-lapse video characteristics in terms of physics, persistence, and variation.
Dataset Type Strong Physics High Persistence Variation
Sky Time-Lapse ${}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{[ACMMM'21]}}}$ (SkyTimelapse, ) General
Time-Lapse-D ${}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{[CVPR'18]}}}$ (TimeLapse-D, ) General
ChronoMagic Metamorphic
Further, we meticulously collect a dataset, ChronoMagic, sourced from Internet time-lapse videos for training our method. As shown in Table 1, our ChronoMagic dataset is distinct from previous time-lapse datasets (SkyTimelapse, ; TimeLapse-D, ), as it consists of metamorphic time-lapse videos (e.g. ice melting, and flower blooming) that show strong physics, high persistence, and variation. Given the domain discrepancy between metamorphic and general videos, we propose the integration of Cascade Preprocessing and Multi-View Text Fusion to capture dynamics into detailed video captions. Contributions are summarized as follows:
  • We introduce MagicTime, a metamorphic time-lapse video generation diffusion model by adopting a standard T2V model with a MagicAdapter scheme. This addresses the limitations of existing methods that are unable to generate metamorphic videos.
  • We design a Dynamic Frames Extraction strategy to effectively extract physics from metamorphic videos and a Magic Text-Encoder to enhance the model’s understanding of metamorphic text prompts.
  • We propose an automatic metamorphic video captioning annotation framework. Utilizing this framework, we have curated a high-quality dataset named ChronoMigic, consisting of 2,265 time-lapse videos, each accompanied by a detailed caption.
  • Extensive experiments demonstrate that our MagicTime is capable of generating high-quality and consistent metamorphic time-lapse videos.
2 Related Work
Text-to-Image Generation    Text-to-image (T2I) synthesis has made significant strides thanks to advancements in deep learning and the availability of large image-text datasets such as CC3M (CC3M, ) and LAION (LAION, ). Compared to variational autoencoders (VAEs) (VAE, ; vqVAE, ), flow-based models (Rezende_Mohamed_2015, ; Dinh_Sohl-Dickstein_Bengio_2016, ), and generative adversarial networks (GANs) (tedigan, ; StyleGAN-XL, ; StyleGAN-T, ), diffusion models (Imagen, ; DALLE, ; DALLE2, ; sdxl, ; lian2023llm, ; Multidiffusion, ) have demonstrated exceptional efficacy in image modeling. For instance, contemporary T2I models often employ Latent Diffusion Models (LDM) (LDM, ), which utilize sampling in latent spaces from pre-trained autoencoders to minimize computational overhead. Recapti (Recapti, ) harnessed the power of Large Language Models (LLM) to improve the text comprehension capabilities of LDM, enabling training-free T2I generation. Furthermore, the integration of additional control signals (ucontrolnet, ; chen2024training, ; t2iadapter, ; dragdiffusion, ; Gligen, ) has further expanded the applicability of Diffusion Models in T2I synthesis. Unlike image generation, the task of video generation required by MagicTime is more complex, as it necessitates both diversity in content across space and consistency in content across time.
Figure 2: The pipeline for high-quality time-lapse video data annotation. (a) Cascade Preprocessing. We successfully detect many useless cuts. The results are obtained by OpenCV. (b) Multi-view Text Fusion. Compared with only inputting keyframes, our pipeline can obtain vivid, detailed and accurate descriptions.
Text-to-Video Generation    The advent of large-scale video-text datasets (webvid, ; Internvid, ; HD-VG-130M, ) has facilitated significant advancements in pre-trained video diffusion models (Make-a-video, ; SVD, ; Align_your_latents, ; ge2023preserve, ; Modelscope, ; Show-1, ). MagicVideo2 (MagicVideo-V2, ) utilized cascaded diffusion to separate image and video generation, starting with text to initially generate images and subsequently produce high-quality videos. Videopoet (Videopoet, ) employed a Transformer-based diffusion approach, combined with multimodal inputs, for zero-shot video generation. Animatediff (animatediff, ) and Lumiere (lumiere, ) extended the 2D-UNet to a 3D-UNet by adding temporal layers and implementing a latent denoising process to conserve computational resources. While these models have achieved basic video generation, the resulting videos typically feature a single scene and a single action, and lack encoding of physical world knowledge. In contrast, our MagicTime incorporates more physics knowledge and offers the capability to generate metamorphic videos.
Time-lapse Video Generation    Previous studies (nam2019end, ; FGLA, ; DtvNet, ; AL, ) on time-lapse video generation predominantly concentrated on general videos, failing to produce the metamorphic videos that we address. For example, (FGLA, ; DtvNet, ) treated time-lapse videos merely as a distinct domain, overlooking the physical knowledge embedded within them. Consequently, the generated videos were limited to scenes with relatively minor movements, such as drifting clouds or moving sunsets. In contrast, our work primarily focuses on generating metamorphic time-lapse videos, which exhibit larger variations and depict the entire process of object metamorphosis. In this work, we aim to incorporate more physical knowledge into MagicTime for metamorphic time-lapse video generation.
3 Methodology
In this section, we commence with a brief overview of the latent diffusion model. Subsequently, we describe the construction of the ChronoMagic dataset. Lastly, we present an outline of MagicTime (refer to Fig. 3) and elaborate on its implementation details.
3.1 Preliminaries
Diffusion models (DDPM, ; DDIM, ) typically consist of two processes: forward diffusion and reverse denoising. Given an input signal $x_{0}$, the process of forward diffusion is defined as follows:
$p_{\theta}\left(x_{t}\mid x_{t-1}\right)=\mathcal{N}\left(x_{t};\sqrt{1-\beta_{t-1}}x_{t-1},\beta_{t}I\right),$ (1)
where $t=1,...,T$ are time steps, and $\beta_{t}\in(0,1)$ is the noise schedule. When the total time steps $T$ are sufficiently large, the resulting $x_{t}$ eventually approximates Gaussian noise $\in N(0,I)$. The purpose of the denoising process is to learn how to reverse the forward diffusion and gradually remove the noise as follows:
$p_{\theta}\left(x_{t-1}\mid x_{t}\right)=\mathcal{N}\left(x_{t-1};\mu_{\theta}\left(x_{t},t\right),\Sigma_{\theta}\left(x_{t},t\right)\right),$ (2)
where $\Sigma_{\theta}(x_{t},t)$ often does not need to be trained but is calculated based on the time step $t$ as a fixed variance. It only needs to predict the mean $\mu_{\theta}(x_{t},t)$ in the reverse process, and this step can also be simplified to training the denoising model $\epsilon_{\theta}(x_{t},t)$ to predict the noise $\epsilon$ of $x_{t}$:
$\mathcal{L}=\mathbb{E}_{,y,\epsilon\sim\mathcal{N}(0,I),t}\left[\left\|\epsilon-\epsilon_{\theta}\left(x_{t},t,\tau_{\theta}(y)\right)\right\|_{2}^{2}\right],$ (3)
where $y$ is text condition, and $\tau_{\theta}(\cdot)$ is the text encoder. By replacing $x_{0}$ with $\mathcal{E}\left(x_{0}\right)$, the latent diffusion model is derived, which also serves as the loss function for MagicTime. In video diffusion models, the parameter $\theta$ typically employs either a Transformer (Videopoet, ; Latte, ) or a U-Net architecture (VideoComposer, ; Modelscope, ; Align_your_latents, ; animatediff, ; zhang2024continuous, ). Considering that the most effective open-source T2V models currently utilize U-Net architecture, we have embraced this architecture in this work.
Figure 3: Overview of the proposed MagicTime approach. We first train MagicAdapter-S to remove the influence of watermarks. Next, MagicAdapter-T is trained to generate metamorphic videos with the help of Dynamic Frames Extraction. Finally, we train a Magic Text-Encoder to enhance text comprehension ability. During the inference stage, all components need to be loaded simultaneously. Slash padding indicates the module is not used.
3.2 ChronoMagic Dataset
Data Curation and Filter    We found that videos in large-scale text-video datasets like WebVid (webvid, ) typically consist of coherent segments depicting everyday life (general videos), devoid of significant changes or object metamorphosis. Consequently, open-source models (SVD, ; Make-a-video, ; animatediff, ) are limited to generating general videos and are unable to produce metamorphic videos. To overcome this limitation, we introduce the ChronoMagic dataset, which includes 2,265 high-quality time-lapse videos with accompanying text descriptions from open-domain content. Time-lapse videos capture complete object metamorphosis and convey more physical knowledge than general videos, thereby enhancing generative models’ understanding of the world.
Given that video platforms host a multitude of low-quality videos, indiscriminate collection would inevitably compromise the dataset’s quality. To mitigate this, we leverage the metadata from video websites for collection and filtering. Our data collection process begins with retrieving original videos from YouTube using "time-lapse" as a search term. Subsequently, we implement a filtering mechanism that excludes videos with short titles, low view counts, or absent hashtags. Furthermore, we discard videos with irrelevant hashtags such as "YouTube," "video," and "shorts." Building on the annotations generated in Section 3.2, we further refine our selection by eliminating irrelevant videos, thus establishing a closed-loop data collection and iteration process. Ultimately, we curate a collection of 2,265 time-lapse videos that meet our specified criteria.
Cascade Preprocessing    Due to complex scene transitions in original videos (see Fig. 2a) affecting the model’s learning of temporal relevance, we need to find a way to solve this problem. The most direct method is to use existing tools for detection, so we choose OpenCV (opencv, ) as a tool. In the initial stage, we define $F$ as the set of all video frames. Each frame $f_{i}$ in $F$ is converted to grayscale $g_{i}\in(1,H,W)$ to mitigate the influence of color on detection. Subsequently, OpenCV is used to calculate the pixel intensity difference $\mathcal{D}_{i,i+1}$ between consecutive frames, yielding the average pixel intensity $\mathcal{S}_{i,i+1}^{opencv}$:
$\mathcal{D}_{i,i+1}=\left|g_{i}-g_{i+1}\right|,\quad\mathrm{for~{}}i=1\mathrm{~{}to~{}}F-1$ (4)
$\mathcal{S}_{i,i+1}^{opencv}=\sum_{h=1}^{H}\sum_{w=1}^{W}\mathcal{D}_{i,i+1}(h,w),\quad\mathrm{for~{}}i=1\mathrm{~{}to~{}}F-1$ (5)
if $\mathcal{S}_{i,i+1}^{opencv}>\theta$, then this position is defined as the transition point. However, we noted that OpenCV is prone to occasional errors. Consequently, we further employ CLIP (CLIP, ) to detect differences between frames:
$\mathcal{S}_{i,i+1}^{clip}=\text{cos}(\operatorname{CLIP}(f_{i}),\operatorname{CLIP}(f_{i+1})),\quad\mathrm{for~{}}i=1\mathrm{~{}to~{}}F-1$ (6)
if $\mathcal{S}_{i,i+1}^{clip}<\vartheta$, then this position is defined as the transition point. The $\text{cos}(\cdot)$ returns a similarity vector. After that, we use a voting mechanism to decide whether a specific location should be recognized as a transition point:
$\mathcal{T}_{i}=\begin{cases}\text{True,}&\text{if }\mathcal{S}_{i,i+1}^{opencv}>\theta\text{ and }\mathcal{S}_{i,i+1}^{clip}<\vartheta\\ \text{False,}&\text{otherwise}\end{cases}$ (7)
where $i$ represents the position of the frame. Ultimately, we segment the video into distinct sections using the identified transition points. Due to possible detection errors in OpenCV and CLIP, the video may still contain transitions that are challenging to identify after processing. In the subsequent stage, we enhance video quality by manually correcting incorrectly cropped segments and then screening clean segments. As illustrated in Fig. 2a, the process results in a high-quality video dataset with fewer transitions.
Multi-view Text Fusion    After preprocessing, it is essential to annotate the video with appropriate captions. A straightforward approach is to input keyframes into a multimodal model for video description. However, the absence of substantial intermediate information may compromise the quality of the annotations. The ChronoMagic dataset’s language modality includes multiview text, such as titles, hashtags, keyframe captions, and video captions, so we leverage these to enhance labeling accuracy. Fig. 2b shows the detailed text generation and fusion process. In videos, titles underscore main themes and actions, while hashtags highlight focus and dynamics. Theoretically, high-quality captions could be generated solely from titles and hashtags. Nevertheless, exaggerated titles and hashtags on the internet can diminish caption quality. To counteract this, we employ in-context Learning and Chain-of-Thought (ICL, ) to guide GPT-4V (gpt4, ; lyu2023gpt, ) in generating keyframe captions using titles and hashtags, thereby reducing model bias and capturing spatial video information. Moreover, temporal video information is crucial. Thus, we use GPT-4V to derive a comprehensive representation of the entire video from keyframe captions, resulting in the final video captions. Additionally, at this stage, GPT-4V evaluates whether the video is a time-lapse and then forms a closed data loop as described in Section 3.2.
3.3 MagicTime Model
MagicAdapter    To transform the pre-trained T2V model to generate metamorphic videos, we decouple the training process into two stages: spatial training and temporal training, refer to Fig. 3.
Internet video data often contains non-removable watermarks, introducing irrelevant noise and data bias. Video generation models consist of spatial and temporal layers—the former focuses on learning visually relevant content, while the latter builds inter-frame continuity. If we can separate temporal and spatial training, allowing the spatial module to concentrate more on the subject rather than the watermark, this will facilitate the temporal layer in learning the motion patterns of metamorphic videos. To accomplish this, in the initial training phase, we remove the temporal layer from the pre-trained model and integrate the MagicAdapter-S into the spatial layer, inspired by QLoRA Adapter (Qlora, ). Subsequently, we freeze the remaining model parameters and employ keyframe-text pairs from the ChronoMagic dataset for training. Formally, let input $X$ to a single spatial layer $\mathcal{F}_{sap}$ and MagicAdapter-S layer $\mathcal{R}_{ada}$ be in the form $\mathbb{R}^{\text{B}\times\text{C}\times\text{H}\times\text{W}}$. The output of this layer is:
$\mathcal{F}_{out}(X)=\mathcal{F}_{sap}(X)+\alpha\times\mathcal{R}_{ada}(X),$ (8)
where $\mathcal{F}_{out}\in\mathbb{R}^{\text{B}\times\text{C}\times\text{H}\times\text{W}}$, and $\alpha$ is a scaling coefficient used to regulate the scope of the adapter. $B$, $C$, $F$, $H$, and $W$ respectively represent the batch size, channel, number of frames, height, and width of the video. Ensuring high temporal consistency is crucial for video generation models. While general videos require consistency within localized segments, metamorphic videos necessitate consistency throughout the entire video, capturing a completely metamorphic process. Training alongside general videos could hinder the model’s comprehension of the new world phenomena. To mitigate this, we separate the training of general and metamorphic videos, thereby maintaining the model’s ability to generate general videos. Specifically, in the second training stage, we reintroduce the pre-trained temporal layer from the first stage, freeze all parameters, and propose a MagicAdapter-T into the temporal layer. Subsequently, we freeze other parameters and conduct training using video-text pairs from the ChronoMagic dataset. If the input $X$ to a single temporal layer $\mathcal{F}_{temp}$ and MagicAdapter-T layer $\mathcal{M}_{ada}$ is in the form of $\mathbb{R}^{\text{B}\times\text{C}\times\text{F}\times\text{H}\times\text{W}}$, then the output of this layer is given by:
$\mathcal{F}_{out}(X)=\mathcal{F}_{temp}(X)+\beta\times\mathcal{M}_{ada}(X),$ (9)
where $\mathcal{F}_{out}\in\mathbb{R}^{\text{B}\times\text{C}\times\text{F}\times\text{H}\times\text{W}}$, $\beta$ is a scaling coefficient, and $B$, $C$, $F$, $H$, and $W$ respectively represent the batch size, channel, number of frames, height, and width of the video. Based on large-scale video-text data (webvid, ), the pre-trained model incorporates general motion capabilities (animatediff, ) through its temporal layer. MagicAdapter-T further enhances the model’s ability to generate metamorphic videos by integrating physical knowledge and predicting metamorphic motion.
Table 2: Comparison of statistics among Sky Time-lapse (SkyTimelapse, ), Time-lapse-D (TimeLapse-D, ), and the proposed metamorphic ChronoMagic dataset. ChronoMagic is the only high-quality dataset for text-to-video generation tasks.
ChronoMagic Sky Time-lapse (SkyTimelapse, ) Time-lapse-D (TimeLapse-D, )
Type Video-Text Video Video
Clips 2,265 2,049 16,874
Resolution 1280 $\times$ 720 640 $\times$ 360 1920 $\times$ 1080
Dynamic Frames Extraction    Open-source algorithms, whether for video generation (360DVD, ; Magic-Me, ) or multimodal models (Video-llava, ; Languagebind, ), typically sample a random continuous sequence of $N$ frames for training. However, this approach captures only a brief segment of the full video, leading to limitations such as monotonous actions and restricted variation in amplitude. This is not conducive to extracting training frames for metamorphic videos. By analyzing ChronoMagic’s characteristics, we opt to sample $N$ frames evenly from a video, capitalizing on the nature of metamorphic videos that encompass the entire process of object metamorphosis. This ensures that the training data exhibit metamorphic properties, rather than being confined to general ones, thereby enhancing the generative models’ ability to comprehend physical phenomena. Nonetheless, uniform sampling from all videos might compromise the model’s original capabilities, resulting in generated videos that lack the action continuity of general videos. Consequently, we retain a portion of general videos and dynamically select the sampling strategy based on the results of cascade preprocessing:
$\mathcal{S}^{*}=\begin{cases}\text{Uniform, with }\mathcal{P}\text{, if }\sum_{i=1}^{F-1}\mathcal{T}_{i}>\delta\\ \text{Random, with }\mathcal{P}\text{, if }\sum_{i=1}^{F-1}\mathcal{T}_{i}\leq\delta\end{cases}$ (10)
where probability $\mathcal{P}$ represents the likelihood of executing the corresponding sampling strategy, enhancing model robustness. $\delta$ represents the threshold for judging transition.
Magic Text-Encoder    Textual descriptions corresponding to metamorphic videos typically contain more temporal and state information than those for general videos. Consequently, using a text encoder trained exclusively on general videos is deemed insufficient for this purpose. To overcome this limitation, we introduce the Magic Text-Encoder, which is specifically designed to encode metamorphic prompts while retaining its ability to process general prompts. We select CLIP as our initial text encoder and incorporate a Low-rank Adapter to encode metamorphic information. The training of the Magic Text-Encoder must be conducted after the overall network training is completed. Formally, given a projection $\mathbf{X}\mathbf{W}=\mathbf{Y}$ with $\mathbf{X}\in\mathbb{R}^{q\times w}$, $\mathbf{W}\in\mathbb{R}^{w\times r}$, adapter computes:
$\mathbf{Y}=\mathbf{X}\mathbf{W}+\gamma\mathbf{X}\mathbf{L}_{1}\mathbf{L}_{2},$ (11)
where $\mathbf{X}$ is the input, $\mathbf{Y}$ is the output, $\mathbf{L}_{1}\in\mathbb{R}^{w\times e}$ and $\mathbf{L}_{2}\in\mathbb{R}^{e\times r}$ are adapter parameters, $\mathbf{W}$ is CLIP parameters, $\gamma$ is a scaling coefficient, and $q$, $w$, $e$, $r$ are the dimensions of input, intermediate, low-rank, and output respectively. To avoid losing original capabilities, only $\mathbf{L}_{1}$ and $\mathbf{L}_{2}$ are updated.
4 Experiment
4.1 Experimental Settings
Figure 4: Qualitative comparison for metamorphic video generation. Among MagicTime, MakeLongVideo (Make-Long-Video, ), ModelScopeT2v (Modelscope, ), VideoCrafter (Videocrafter, ), Zeroscope (zeroscope, ), LaVie (lavie, ), T2V-Zero (T2V-zero, ), Latte (Latte, ) and Animatediff (animatediff, ), our MagicTime achieves the best results among all T2V models.
Datasets    Metamorphic video generation represents a novel paradigm, yet dedicated evaluation datasets are scarce, presenting challenges for thorough method assessment. To address this need, we utilize the ChronoMagic dataset, the first benchmark specifically designed for metamorphic video generation. Detailed information and examples are provided in the Appendix. This dataset consists of high-quality metamorphic time-lapse videos with corresponding captions. The statistics of the ChronoMagic dataset are presented in Table 2.
Metrics    In this work, we utilize three metrics for quantitative evaluation: Fréchet Inception Distance (FID) (FID, ), Fréchet Video Distance (FVD) (FVD, ) and CLIP Similarity (CLIPSIM) (CLIPSIM, ). These are standard indicators assessing generated image quality, video quality, and text-content similarity. For a fair comparison, we set fixed random seeds in all experiments.
Figure 5: Examples of our approach applied to generated metamorphic videos. Metamorphic videos categories include RealisticVision (Row 1-3), RcnzCartoon (Row 4-6) and ToonYou (Row 7-9). Even with a large state amplitude, the quality, consistency, and semantic relevance of generated videos are excellent.
Implementation Details    We choose Stable Diffusion v1.5 (LDM, ) and Motion Module v15 (animatediff, ) as our baseline models. In the training stage, the resolution is set to 512$\times$512, 16 frames are taken by Dynamic Frames Extraction from each video, the batch size is set to 1,024, the learning rate is set to 1e-4, the total number of training steps is 10k, and the classify free guidance random null text ratio is 0.1. We use a linear beta plan as the optimizer with $\beta_{\text{start}}=8.5e-4$ and $\beta_{\text{end}}=1.2e-2$. We set $\theta$, $\vartheta$, $\mathcal{P}$ and $\delta$ in Sections 3.2 and 2 to 40, 0.5, 0.9 and 3 respectively to enhance robustness. In the inference stage, we utilize DDIM with 25 sampling steps, a text-guided scale of 8.0, and a sampling size of 512$\times$512. We have collected personalized SD models from CivitAI to verify the generalizability of our method, including RealisticVision, RcnzCartoon, and ToonYou. Due to space limitations, more qualitative experiments on generated metamorphic videos are shown in the Appendix. Further, selected examples from our ChronoMagic dataset, along with the questionnaire employed for human evaluation, are provided in the Appendix for further reference.
Table 3: Quantitative comparison with state-of-the-art T2V generation methods for the text-to-video task. Results are produced on ChronoMagic. Our method achieves either the best results or second-best results across all metrics. "normal-$\downarrow$" denotes lower is better. "normal-$\uparrow$" denotes higher is better.
Method Venue Backbone FIDnormal-$\downarrow$ FVDnormal-$\downarrow$ CLIPSIMnormal-$\uparrow$
MakeLongVideo (Make-Long-Video, ) Github’23 U-Net 143.07 511.62 0.2852
ModelScopeT2V (Modelscope, ) Arxiv’23 U-Net 61.49 500.43 0.3093
VideoCrafter (Videocrafter, ) CVPR’23 U-Net 54.61 473.69 0.3149
Zeroscope (zeroscope, ) CVPR’23 U-Net 103.74 497.60 0.3079
LaVie (lavie, ) Arxiv’23 U-Net 62.28 445.60 0.3090
T2V-zero (T2V-zero, ) ICCV’23 U-Net 58.58 488.85 0.3203
Latte (Latte, ) Arxiv’24 DiT 59.10 445.85 0.2939
Animatediff (animatediff, ) ICLR’24 U-Net 63.83 451.35 0.3156
MagicTime Ours U-Net 58.12 441.17 0.3272
Figure 6: (a) Human evaluation between MagicTime and state-of-the-art T2V generation methods. MagicTime is preferred by human voters over other methods. (b) Human evaluation of different ablation settings. The most crucial components are MagicAdapter-S/T (MA-S/T) and Dynamic Frames Extraction (DFE), which empower MagicTime with metamorphic video generation capabilities. Magic Text-Encoder (MTE) and Multi-view Text Fusion (MTF) are primarily used to enhance video quality.
4.2 Qualitative Analysis
In this section, we first compare our method, MagicTime, with several state-of-the-art models including MakeLongVideo (Make-Long-Video, ), ModelScopeT2V (Modelscope, ), VideoCrafter (Videocrafter, ), Zeroscope (zeroscope, ), LaVie (lavie, ), T2V-zero (T2V-zero, ), Latte (Latte, ), and Animatediff (animatediff, ) in the task of metamorphic video generation. All input text is sourced from our ChronoMagic dataset. The evaluated tasks involve generating videos that depict significant metamorphosis, such as the construction of a modern house in Minecraft, the blooming of pink plum blossoms, the baking of bread rolls, and the melting of ice cubes, as shown in Fig. 4. Our evaluation focuses on Frame Consistency (FC), Metamorphic Amplitude (MA), and Text Alignment (TA). While all methods, except MakeLongVideo, produce high-quality videos with strong temporal modeling implied high FC, they struggle with TA and MA. The limited physical priors in existing text-to-video models restrict them to general scene generation, such as swaying flowers (ModelScopeT2V, VideoCrafter, Latte), static buildings (Zeroscope, LaVie, Animatediff, T2V-zero), which are incorrect contents. In contrast, MagicTime excels in capturing metamorphic transitions smoothly and accurately reflecting physical sequences, such as layer-by-layer Minecraft construction and natural flower blooming. Notably, MagicTime demonstrates the ability to generate visually consistent content over large time spans, accurately depicting processes like ice melting and dough baking, further validating the model’s effectiveness.
Next, to demonstrate the robustness and generality of our approach, we present several text-guided metamorphic video generation results. As shown in Fig. 5, MagicTime effectively converts a general T2V model into a metamorphic video generator, yielding impressive outcomes ranging from swelling and melting to growing and building. These results showcase a robust understanding of physical phenomena encoded within the model. Even when the textual prompt requires the generation of content spanning long periods, the generated videos maintain high quality, consistency, and semantic relevance. This success is attributed to our method of preserving image generation and temporal modeling priors learned from SD and T2V baselines on large datasets while encoding physical priors characteristic of metamorphic videos.
Figure 7: Ablation study for different parts of MagicTime. It shows the effectiveness of Magic Text-Encoder (MTE), Multi-view Text Fusion (MTF), MagicAdapter-S/T (MA-S/T), and Dynamic Frames Extraction (DFE).
4.3 Quantitative Analysis
In this section, we conduct a comprehensive quantitative evaluation to assess the performance of our approach. We compare the generation quality on the ChronoMagic dataset between MagicTime and baseline methods, as presented in Table 3. Consistent with the results in Fig. 4, MakeLongVideo exhibits the highest FID and FVD values, indicating its inferior visual quality compared to other methods. While VideoCrafter has the lowest FID, our method boasts the lowest FVD, suggesting that MagicTime offers the highest video quality. CLIPSIM, originally applied in image generation, mainly measures the similarity between individual video frames and text for videos, which limits its effectiveness in assessing video generation quality. Given the lack of superior metrics in the T2V field, we use CLIPSIM as an indicator of video-text relevance. Numerically, although the video-text relevance of MagicTime is the highest, it is only slightly higher than the other methods. However, as shown in Fig. 4, only MagicTime adheres to the text prompt instructions to generate metamorphic videos, while the baseline methods generate general videos based on the theme of the prompt, implying that their CLIPSIM scores should be lower. Overall, our method surpasses all compared methods, confirming that in modeling metamorphic videos, we can ensure superior video generation quality while successfully incorporating more physical priors, thus producing results that align with the definition of metamorphic videos.
4.4 Human Evaluation
As previously mentioned, metrics such as FID, FVD, and CLIPSIM may not fully capture the generative capabilities of T2V methods. To address this, we enlisted 200 participants to evaluate the Visual Quality, Frame Consistency, Metamorphic Amplitude, and Text Alignment of videos generated by MagicTime and baseline methods. We selected 15 sets of samples for each criterion and asked participants to choose the high-quality metamorphic video that best aligned with their expectations. For fairness, all options were randomized, and each participant could only complete one questionnaire under identical conditions. Due to the need for a large pool of participants in the evaluation, we only select the top 5 T2V methods for voting. We finally collected 185 valid questionnaires. The results, depicted in Fig. 5(a), demonstrate that our method significantly outperforms other leading T2V methods. Furthermore, MagicTime significantly enhances the quality of metamorphic video generation, validating our approach’s effectiveness in encoding physical knowledge through training with metamorphic videos. Visualizations of the questionnaire are provided in the Appendix.
Figure 8: Effect of different large multimodal models and annotation strategies for the generated captions. Gray indicates inaccuracy or redundancy descriptions, red indicates precise or detailed descriptions. Text prompts exhibit clear temporal details, abundant object details, and smooth narrative flow when adopting GPT-4V and Multi-view Text Fusion.
Figure 9: Effect of the number of inversion steps $t$. Best results are obtained when $t=25$.
4.5 Ablation Study
Validation on Different Components    We primarily conduct ablation studies on our proposed strategies of Dynamic Frames Extraction (DFE), MagicAdapter-T (MA-T), MagicAdapter-S (MA-S), Multi-view Text Fusion (MTF) and Magic Text-Encoder (MTE), as illustrated in Fig. 7. Given the prompt "A bud transforms into a yellow flower", the Group w/o MTE, can produce relatively better results. However, compared to the complete MagicTime, it still slightly lags in aspects such as fluency and degree of metamorphosis. The Group w/o MTF, due to the use of low-quality prompts, resulted in a decrease in generation quality, such as the appearance of strange buds in the middle of flowers. The Group w/o MA-S can generate the complete process, but as the dataset contains a large amount of watermark data, the results also include watermarks. The Group w/o MA-T struggles to effectively encode the concept of metamorphic, resulting in general video without metamorphic characteristics. The Group w/o DFE, similarly to the Group w/o MA-T, can not generate metamorphic video and has reduced generation quality due to the lack of metamorphic feature acquisition. We also conduct quantitative experiments, as shown in Fig. 5(b), demonstrating that MagicTime significantly outperforms other methods across all four dimensions. The experimental settings are consistent with those in Section 4.4, and we finally collected 115 valid questionnaires. Our MagicTime significantly enhances metamorphic video generation. With simple adapters and training strategies, it validates the hypothesis that incorporating time-lapse videos into the model can encode more physical knowledge.
Effect of the Enhanced Text Prompt    There are various methods for video captioning annotation. For example, 360DVD (360DVD, ) employs BLIP (blip, ), which is tailored for captioning tasks, while LanguageBind (Languagebind, ) uses the multimodal large model OFA (OFA, ). However, we observe that both specialized and general models have limitations in comprehending metamorphic videos. We illustrate this by examining the leading multimodal large model Video-LLaVA (Video-llava, ), with results in Fig. 8. It shows Video-LLaVA struggles to capture fine-grained details and often produces irrelevant descriptions that may interfere with prompt instructions. Therefore, we choose GPT-4V for video annotation. Nevertheless, since GPT-4V only processes keyframes, it risks losing valuable information in the entire video. To mitigate this, we introduce Multi-view Text Fusion (MTF) to enhance the quality of video captions. GPT-4V with MTF can generate text prompts that are rich in temporal details, object specifics, and narrative coherence. We also apply MTF to Video-LLaVA; however, due to its limited ability to follow instructions, the improvements pre- and post-application are relatively minor.
Effect of the Number of Inversion Steps    To assess the influence of the number of inversion steps, we conducted an ablation study to examine the effect of varying $t$ (i.e., the number of inversion steps) during the inference stage of MagicTime. We set $t=5,15,25,35$ and applied our approach to the ChronoMagic dataset to obtain metamorphic results. We evaluated Visual Quality (VQ), Frame Consistency (FC), Metamorphic Amplitude (MA), and Text Alignment (TA) for each $t$ value, with examples illustrated in Fig. 9 using RcnzCartoon as a representative case. In terms of VQ, we observed that when $t$ is less than 25, the larger $t$ is, the better VQ becomes. For FC, after $t$ exceeds 5, there is good consistency throughout; for MA, when $t$ equals 25, even if it is increased further, MA only shows minor changes; for TA, regardless of the value of $t$, it can always generate results that are aligned with a text prompt. When $t$ is too small, the diffusion process lacks the necessary flexibility for significant changes, challenging effective generation. Conversely, a larger $t$ value leads to extended reasoning times. In practice, we set $t=25$ to achieve best performance.
4.6 Integrate MagicTime Scheme into DiT-based Architecture
Sora videoworldsimulators2024 is a text-to-video generation model released by OpenAI in February 2024, which shows the powerful potential of simulating the physical world. Due to OpenAI’s closed-source policy, the academic community is currently working on reproducing the open-source Sora-like structure based on the DiT (Diffusion-Transformer)-based Text-to-Video (T2V) generation framework Latte ; opensora ; tinysora . The mission of this project is to help reproduce Sora and provide high-quality video-text data and data annotation pipelines, to support Open-Sora-Plan or other DiT-based T2V models. To this end, we take an initial step to integrate our MagicTime scheme into the DiT-based Framework. Specifically, our method supports the Open-Sora-Plan v1.0.0 (tinysora, ) for fine-tuning. The Open-Sora-Plan v1.0.0 exhibits superior performance in generating landscape videos. Therefore, we first scale up with additional metamorphic landscape time-lapse videos in the same annotation framework to get the ChronoMagic-Landscape dataset. Then, we fine-tune the Open-Sora-Plan v1.0.0 with the ChronoMagic-Landscape dataset to get the MagicTime-DiT model. As shown in Fig. 10, our MagicTime-DiT has generated high-quality metamorphic landscape time-lapse videos. Now, we are still exploring to support more DiT-based T2V structures and continue on the road to scaling our ChronoMagic dataset, in hopes of truly bringing useful insights to the community.
Figure 10: Examples of metamorphic time-lapse videos generated by our MagicTime-DiT based on Open-Sora-Plan v1.0.0 tinysora with ChronoMagic-Landscape Dataset.
5 Discussion and Conclusion
In this paper, we present MagicTime, a novel framework for metamorphic time-lapse video generation. It can generate metamorphic videos conforming to the content distribution and motion patterns in real-captured metamorphic videos. Extensive experiments demonstrate our effectiveness in generating high-quality metamorphic videos with various prompts and styles. Further, our MagicTime model and ChronoMagic benchmark can provide a simple yet effective solution for metamorphic video generation, which can facilitate further investigation by the community.
Ethics Statement.    MagicTime can generate not only metamorphic videos with large time spans and high levels of realism but also high-quality general videos. Potential negative impacts exist as the technique can be used to generate fake video content for fraudulent purposes. We note that any technology can be misused.
Appendix A Additional Experimental Results
A.1 Pseudo Long Video Generation
Current text-to-video (T2V) models [40, 3, 1, 54] often face challenges in generating 16-frame videos, which typically feature only camera motion. The generation of longer videos [71] commonly involves multiple forward inferences based on a 16-frame model, followed by concatenating the results. Since our method can produce metamorphic videos within 16 frames, it effectively compresses the content of a longer video into 16 frames. Consequently, combining a frame interpolation model with the outputs of MagicTime can theoretically achieve the effects of longer video generation. We hope this insight will inspire and advance future research in this burgeoning field. We present some examples using the interpolation model [72], as illustrated in Fig. 12. It is evident that with MagicTime’s capability for metamorphic generation, pseudo-long videos can be produced.
Figure 11: Ablation on the settings of $\theta$. Best results are obtained when $\theta=40$.
A.2 Resolution Scalability
Resolution scalability is a significant attribute of the T2V model with a U-Net architecture. To assess whether MagicTime retains this feature, we conducted related experiments, as depicted in Fig. 13. Despite being trained at a resolution of $512\times 512$, our method demonstrates remarkable generalization capabilities across a wide range of resolutions. This scalability is particularly prominent when our model is applied to higher resolutions such as $1024\times 1024$, where it maintains high fidelity and consistency in generated details. Similarly, at lower resolutions, such as $128\times 128$ and $256\times 256$, our method preserves structural integrity and aesthetic quality. This robust generalization may be attributed to the inherent properties of the underlying generative model, Stable Diffusion [25], which we employ in our framework.
Figure 12: Pseudo Long Video Generation. Our method can be converted to a long video generation model by using an interpolation model [72].
Figure 13: Resolution Scalability. The results demonstrate the generalization of our method.
A.3 More Cases on Metamorphic Videos
Due to space constraints, we present additional metamorphic videos generated by MagicTime and other baseline methods in Fig. 14 and Fig. 16. It is clear that MagicTime consistently surpasses existing state-of-the-art Time-to-Video (T2V) models across various scenarios, from melting ice to the expansion of bread. This highlights MagicTime’s adeptness in capturing physical phenomena such as expansion, melting, and burning. Furthermore, our approach consistently yields expected results and demonstrates robust generalization capabilities across diverse DreamBooths, from RealisticVison and ToonYou to RcnzCartoon. Fig. 17 showcases samples from the dataset, which primarily consists of metamorphic time-lapse videos of plants, buildings, ice, food, and other categories.
A.4 More Hyperparameter Analysis
In the Cascade Preprocessing module, we employ OpenCV to detect video transitions based on the grayscale difference between two frames. The choice of the difference threshold $\theta$ significantly impacts OpenCV’s ability to identify transitions. Therefore, we conducted ablation experiments with varying $\theta$ values to find optimal values, as depicted in Fig. 11. The actual average number of transitions is approximately six. When $\theta$ is too large, many frames are erroneously identified as transitions. Conversely, when $\theta$ is too small, it results in missed detections. It can be seen that optimal results are achieved when $\theta$ is set to 40.
Figure 14: More qualitative comparison of metamorphic video generation. MagicTime achieves better results compared to the current leading T2V models.
Figure 15: Visualization of the Questionnaire for Human Evaluation.
Figure 16: More examples of our approach applied to generated metamorphic videos. Metamorphic videos categories include RealisticVision (Row 1-4), RcnzCartoon (Row 5-6) and ToonYou (Row 7-8). Even with a large state amplitude, the quality, consistency, and semantic relevance of generated videos are excellent.
Figure 17: Samples from the ChronoMagic dataset. The dataset consists of metamorphic time-lapse videos, which exhibit physical knowledge, long persistence, and strong variation.
A.5 Questionnaire Visualization
To enhance the reliability of the experiment, we developed both coarse-grained and fine-grained questionnaires during the data collection phase and invited 200 participants to provide their responses. The coarse-grained questionnaire required participants to assess these four criteria and select the video that best met their requirements, as shown in Fig. 15a. As this questionnaire was more user-friendly, a total of 300 questionnaires were collected for all surveys related to this work. The fine-grained questionnaire asked participants to evaluate the visual quality, frame consistency, metamorphic amplitude, and text alignment of each video set, as shown in Fig. 15b.